İf the supply curve is q = 3 + 1.5p, what is the producer surplus if the price is 12?

1
Expert's answer
2021-11-01T19:23:20-0400

q = 3 + 1.5p

p=12

"q=3+1.5\times12=3+18=21"

Let Pmin=4

"producer surplus=0.5(Po-Pmin)\times q=0.5(12-4)\times21=84"